For what values of c are the vectors [βˆ’8, 𝑐, 4] and [𝑐, 𝑐 2 , 𝑐] orthogonal?


1
Expert's answer
2021-10-20T09:56:01-0400

Let "\vec c_1=-8\hat{i}+c\hat{j}+4\hat{k}, \vec c_2=c\hat{i}+c^2\hat{j}+c\hat{k}"

For vectors to be orthogonal, "\vec c_1.\vec c_2=0"

"\therefore-8c+c^3+4c=0\\
\Rightarrow c(c^2-4)=0\\
\Rightarrow c=0,2,-2"